What is the diagonal of a rectangular sheet of paper if you know that's its area is 630cm squared and one side is 30cm long?
1
Expert's answer
2013-03-22T04:22:16-0400
If the area of a paper is A = 630[cm²] and one side is L = 30[cm] long the other side is

W = A/L = 630[cm²]/30[cm] = 21[cm]

long. According to the Pythagoras' theorem, the diagonal of a sheet of paper is

D = √(W² + L²) = √((30[cm])² + (21[cm])²) = √1341[cm] ≈ 36.62[cm]

long.